Write the equation of the line containing the given point and perpendicular to given line. (0,6) 3x+2y=7
Answer by lwsshak3(6761)

Write the equation of the line containing the given point and perpendicular to given line. (0,6) 3x+2y=7
**
standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
for given equation: 3x+2y=7
2y=-3x+7
y=-3x/2+7/2
slope=-3/2
slope of line perpendicular to given line=2/3 (negative reciprocal of -3/2)
Equation of perpendicular line: y=2x/3+b
solving for b using given point on the line (0,6)
3y=2x+3b
3*6=2*0+3b
18=3b
b=6
Equation of perpendicular line:
y=2x/3+6